A group of 85 students in first-year biology at Western University just wrote their second midterm. Because the first year tests involve a group component in which students learn the correct answers to the midterm questions at the end of the exam, they each know what they scored on the test. Based on their individual results, they computed a mean midterm mark of 19 (it was a test marked out of 25), with standard deviation of 4.3 marks. They considered themselves a simple random sample of all first-year biology students, and consequently calculated a confidence interval to estimate the mean midterm score for the test of 17.893 to 20.107 marks). What confidence level was used to create this confidence interval?

Correct answer is 98% i need process work
Answer #1

Here we have :

n = 85

Mean = 19

Standard deviation = 4.3

AzObHt477UdxAAAAAElFTkSuQmCC

(98 % confidence level)

df = n-1 = 85 -1 = 84

Critical value :

tc = t∝/2, df = t0.02/2, 84 = 2.372

Confidence interval :

4.3 ув5 19 + 2.372 x 19 t 1.107 17.893 20.107